OpenStudy (anonymous):

A garden is 14 1/2 ft by 13 1/4 ft. What is the area of the garden? HELPPPPP ME PLEASE!!

OpenStudy (anonymous):

multiply... \(\large 14\frac{1}{2} \times 13\frac{1}{4}\) = ???

OpenStudy (anonymous):

192.1 square ft

OpenStudy (anonymous):

Pathosdebater thank you so much you are helping me on everuthing! :D